How many grams of calcium phosphate are theoretically produced ifwe start with 3.40 moles of...

Consider the balanced equation below. How many grams of calcium phosphate are theoretically produced if we start with 3.40 moles of \(\mathrm{Ca}\left(\mathrm{NO}_{3}\right)_{2}\) and 2.40 moles of \(\mathrm{Li}_{3} \mathrm{PO}_{4} ?\) Hint: Consider limiting reagent.

$$ 3 \mathrm{Ca}\left(\mathrm{NO}_{3}\right)_{2}+2 \mathrm{Li}_{3} \mathrm{PO}_{4} \rightarrow 6 \mathrm{LiNO}_{3}+\mathrm{Ca}_{3}\left(\mathrm{PO}_{4}\right)_{2} $$
Answer #1
First we must get the balanced equation

2.4 mol of ?

I will write the net ionic

3 Ca2+ + 2 PO43- = Ca3(PO4)2

the ratio between Ca2+ and PO43- is 3 : 2

moles PO43- required = 3.40 x 2 /3 = 2.27 so PO43- is the limiting reactant

the ratio between PO43- and Ca3(PO4)2 is 2 : 1

moles Ca3(PO4)2 = 2.27 x 1 /2 = 1.13

mass calcium phosphate = 1.13 mol x 310.18 g/mol=350.5 g
Answer #2
3 Ca(NO3)2 + 2 Li3PO4 >> 6 LiNO3 + Ca3(PO4)2

the ratio between Ca(NO3)2 and Li3PO4 is 3 :2

3 : 2 = 3.40 : x

x = 2.26 moles Li3PO4 needed for the reaction

we have 2.40 moles Li3PO4 so it is in excess

Ca(NO3)2 limiting reactant

The ratio between Ca(NO3)2 and Ca3(PO4)2 is 3 : 1

3 : 1 = 3.4 : x

x = 1.13 moles Ca3(PO4)2 produced

Molecular mass Ca3(PO4)2 = 310.18 g/mol

grams = 310.18 g/mol x 3.40 moles = 10546